OpenStudy (anonymous):

Simplify 27^-(2/3) 9 1/9 -1/9 −9

OpenStudy (luigi0210):

\[\LARGE (\frac{a}{b})^{-x}=(\frac{b}{a})^{x}\]

OpenStudy (anonymous):

so i make the 9 into a frction like \[\frac{ 9 }{ 1 }\] then flip it?

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

ok then what would i do with the -2/1 exponent?

OpenStudy (anonymous):

i meant 2/3

OpenStudy (luigi0210):

When you flip the fraction, it just becomes positive.

OpenStudy (anonymous):

would i multiply it with 9/1?

OpenStudy (luigi0210):

It goes like this: First you flip the number to make the exponent positive \[\LARGE 27^{-2/3}=(\frac{1}{27})^{2/3}\] Now just distribute the exponent out to everything: \[\LARGE a^{2/3}=\sqrt[3]{x^2}\]

OpenStudy (anonymous):

but how would i get it to one of my answer choices

OpenStudy (luigi0210):

So what's 1^(2/3) and 27^(2/3)?

OpenStudy (anonymous):

1^(2/3) is \[^{3}\sqrt{1^{2}}\] and the other is \[^{3}\sqrt{27^{2}}\]

OpenStudy (luigi0210):

Correct, now simplify them

OpenStudy (anonymous):

how would i do that

OpenStudy (luigi0210):

Well 1 to any power 1. So that one is simple. For the 27, first take the cubic root of it.

OpenStudy (anonymous):

so i square it then divide it by 3?im not sure how to do tht

OpenStudy (luigi0210):

Hmm, do you have a calculator?

OpenStudy (anonymous):

yea it is on my computer.i have it up right now

OpenStudy (luigi0210):

Okay, type this in: 27^(1/3)

OpenStudy (anonymous):

i got it.it is 3

OpenStudy (luigi0210):

Since \[\LARGE 27^{2/3}=27^{1/3*(2)}\] Right, now sqaure that and you get..

OpenStudy (luigi0210):

*square

OpenStudy (anonymous):

i get 9

OpenStudy (luigi0210):

Correct! So our final answer is: \[\LARGE \frac{1}{9}\]

OpenStudy (anonymous):

thank you so much i really apreciate it

OpenStudy (luigi0210):

ANytime bud.

OpenStudy (anonymous):

thanks man